[Brahms-l] Shift report 20050228 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Feb 28 2005 - 07:57:05 EST
Shifter leader was HHD
Shifters were 


1. Good beam was on most of the time. There was a minor interruption (~15 min) due to MCR doing gap cleaning. Beam studies will take place 0700-1900. 

2. Runs taken: 13523,13524,13525,15326
   FS@8A2442: 41k, 72k, 83k, 39k         
   MRS@40B700: 248k, 93k, 376k, 237k

3. No problems.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 28 07:57:39 2005

This archive was generated by hypermail 2.1.8 : Mon Feb 28 2005 - 07:57:48 EST